# On an entropy inequality for quadratic forms and applications

### v1

@@ Abstract
Let $\phi(x,y)$ be a non-degenerate rational quadratic form. Let $X$ and $Y$ be independent $(s,C)$-Frostman random variables whose ranges are contained in $[-c_{1},c_{1}]$, with $10<s<$ 1, $C,c_{1}\geq 1$. We prove that there exist a positive constant $\epsilon=\epsilon(s,\phi)$ and an integer $N=N(s,C,c_{1},\phi)$ such that
@@
for all $n>N$. The proof introduces a novel multi‑step entropy framework, combining the submodularity formula, the discretized entropy Balog-Szemerédi-Gowers theorem, and state-of-the-art results on the Falconer distance problem, to reduce general forms to a diagonal core case. As an application, we derive a result on a discretized sum-product type problem. In particular, for a $\delta$-separated set $A\subset[0,1]$ of cardinality $\delta^{-s}$, satisfying some non-concentration conditions, there exists $\epsilon=\epsilon(s,\phi)>0$ such that
$\mathcal{E}_{\delta}(A+A)+\mathcal{E}_{\delta}(\phi(A,A))\gg\delta^{-\epsilon}(\#A)$
for all $\delta$ small enough. Here by $\mathcal{E}_{\delta}(A)$ we mean the $\delta$-covering number of $A$.
@@ 1 Introduction
More precisely, we are interested in bounding the following quantity $\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}$ from below, where $\phi(x,y)=a_{1}x^{2}+a_{2}xy+a_{3}y^{2}$ is a non-degenerate rational quadratic form, i.e. $a_{2}^{2}-4a_{1}a_{3}\neq 0$, $X$ and $Y$ are independent $(s,C)$-Frostman random variables whose ranges are bounded, and $s\in(0,1)$. We state our main result as follows.
@@
Let $\phi(x,y)$ be a non-degenerate rational quadratic form. Let $X$ and $Y$ be independent $(s,C)$-Frostman random variables whose ranges are contained in $[-c_{1},c_{1}]$, with $10<s<$ 1, and $C,c_{1}\geq 1$. Then there exist a positive constant $\epsilon=\epsilon(s,\phi)$ and an integer $N=N(s,C,c_{1},\phi)$ such that for all $n>N,$ we have
$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon).$ (1)
To show the sharpness of this theorem, we summarize the main statements of several examples in Section A:
@@
    Any $(s,C)$-Frostman random variable $Z$ has the property that $H_{n}(Z)\geq ns-\log C+O(1)$ (see Example A.1 for an example matching this bound). The lower bound in (1) says that the discretized entropies of $X+Y$ and $\phi(X,Y)$ cannot be small simultaneously. We show in Examples A.1 and A.2 that at least one of the terms $H_{n}(X+Y)$ or $H_{n}(\phi(X,Y))$ can be as small as $(1+o(1))ns$.
@@
    The positive $\epsilon$ might not exist if one of the conditions of the theorem is not satisfied, including when $s=0$ or $1$ (see Example A.1), the Frostman constant $C$ grows with $n$ (see Example A.4), or the polynomial is degenerate (see Example A.3).
@@
    In the statement of Theorem 1.1, the fact that $N$ depends on $C$ , and the condition that $X$ and $Y$ are independent are both essential (see Examples A.4 and A.5, respectively).
@@
    For any $10<s<$ 1 and any sufficiently large $n\in\mathbb{N}$, there exists an $(s,O(1))$-Frostman random variable $X$ with range contained in $[0,1]$, such that
@@
Let $\phi(x,y)$ be a non-degenerate rational quadratic form. For $10<s<$ 1, there exists a positive constant $\epsilon=\epsilon(s{\color[rgb]{0,0,0},\phi})$ such that for every Borel set $A\subset{\color[rgb]{0,0,0}[0,1]}$ with $H^{s}(A)>0$, we have
$\max\left\{\underline{\dim}_{B}(A+A),\underline{\dim}_{B}(\phi(A,A))\right\}\geq s+\epsilon.$
This theorem is of independent interest, particularly when $\phi(x,y)$ is a non-bilinear form, and might be considered as an application of the Falconer distance problem. If $\phi(x,y)=xy$, then stronger results have been obtained by using a variety of techniques, see [4, 10, 11, 16] and references therein for current progress. Given this line of research, it is reasonable to expect different proofs of Theorem 1.2 with other methods. In the prime field setting, such a result has been proved by using tools from incidence geometry, see [8, Theorem 1.5] and [14, Theorem 1.3], for example.
@@
Let $\phi(x,y)$ be a non-degenerate rational quadratic form. Given $10<s<$ 1 and a constant $C\geq 1$. Then there exist $\epsilon=\epsilon(s{\color[rgb]{0,0,0},\phi})>0$ and $\delta_{0}=\delta_{0}(s,C,\phi)>0$ such that the following holds for all $\delta\leq\delta_{0}$. Let $A\subset{\color[rgb]{0,0,0}[0,1]}$ be a $\delta$-separated set of cardinality $\delta^{-s}$. Suppose that for all intervals $I$of length at least $\delta$, we have the non-concentration estimate
#$\#(A\cap I)\leq C|I|^{s}(\#A).$
@@
$\mathcal{E}_{\delta}(A+A)+\mathcal{E}_{\delta}(\phi(A,A))\gg C^{-O(1)}\delta^{-\epsilon}(\#A).$
We now sketch the main ideas in the proof of Theorem 1.1.
Our base case is corresponding to the diagonal forms $\phi(X,Y)=\alpha X^{2}+\beta Y^{2}$ (Lemma 4.1), with independent $(s,C)$-Frostman random variables $X,Y$. Roughly speaking, by applying the submodularity formula (Lemma 2.18) on certain algebraic relations between the sum function $X+Y$ and the distance function $(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}$, we are able to reduce the problem to bounding the entropy of the distance function from below. In particular, let $W$ be an $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subset\mathbb{R}^{2}$, where $10<s<$ 1, the constant $C\geq 1$ is bounded and does not depend on $n$, and $R>0$. Let $W^{\prime}$ be a random variable which is i.i.d with $W$. Then there exists a constant $\epsilon_{0}=\epsilon_{0}(s)>0$ such that
@@
for all $n\geq N_{0}$, where $N_{0}=N_{0}(s,C,R)$. Note that if the Frostman constant $C$ depends on $n$, then the inequality (2) might not be true (see Example A.5). To prove (2), we make use of results on the Falconer distance problem. While classical estimates taken from [12, Section 4.2] are sufficient in the case $s>1/2$, we invoke the recent work of Shmerkin and Wang [20] for $s\leq 1/2$.
For the general form $\phi(X,Y)$, the proof becomes much more complicated. The first step is to transform it to a diagonal form $\tilde{\phi}(\tilde{X},\tilde{Y})$, but the transformed variables $\tilde{X}$ and $\tilde{Y}$ are no longer independent. To proceed further, we introduce a novel approach, which can be briefly described as follows. By applying the discretized entropy version of Balog-Szemerédi-Gowers theorem (see Theorem 3.1), we obtain new random variables $X^{\prime\prime}$ and $Y^{\prime\prime}$ that are conditionally independent, and thus reduce the problem accordingly. Generally speaking, there exists a random variable $(X^{\prime\prime},Y^{\prime},X^{\prime},Y^{\prime\prime})$ such that
(i) $(X^{\prime},Y^{\prime})\equiv(\tilde{X},\tilde{Y})$;
@@
Large dimension: Let $W$ be an $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subset\mathbb{R}^{2}$, where $11/2<s<$ 1, $C\geq 1$ and $R>0$. Let $W^{\prime}$ be a random variable which is i.i.d with $W$. Then for any number $t$ satisfying $t\leq 1$ and $s<t<2s-1/2$, we have that for all $n\in\mathbb{N},$
@@
Small dimension: Let $X^{\prime\prime}$, $Y^{\prime\prime}$ be independent random variables with ranges contained in $[-c,c]$, with $c>0$. Suppose that $X^{\prime\prime}$ and $Y^{\prime\prime}$ are $(s,C_{1})$-Frostman and $(s,C_{2})$-Frostman, respectively, where $20<s\leq 1/$ 2 and $C_{1},C_{2}\geq 1$. Let $(X^{\prime\prime\prime},Y^{\prime\prime\prime})$ be i.i.d. with $(X^{\prime\prime},Y^{\prime\prime})$. Then for any $t$ with $s<t<\frac{s+\varphi(2s)}{2}$, for all $n\in\mathbb{N}$, we have
@@
Plugging these inequalities in our method, the theorem follows.

### v2

@@ Abstract
We introduce Frostman conditions for bivariate random variables and study discretized entropy sum–product phenomena in both independent and dependent regimes.
In the independent setting, we show that for any non-degenerate rational quadratic form $\phi(x,y)$ and $0<s<1$, there exists some positive constant $\epsilon=\epsilon(\phi,s)$ such that any independent marginally $(s,C;s,C)$-Frostman random variable $(X,Y)$ satisfies
@@
for sufficiently large $n$. The proof introduces a novel multi‑step entropy framework, combining the submodularity formula, the discretized entropy Balog-Szemerédi-Gowers theorem, and state-of-the-art results on the Falconer distance problem, to reduce general forms to a diagonal core case.
In the dependent setting, we obtain analogous bounds for special polynomials. As an application, we derive that there exists $\epsilon=\epsilon(s)>0$ such that, for any $\delta$-separated set $A\subset[0,1]$ of cardinality $\delta^{-s}$ satisfying some non-concentration condition and any subset $G\subseteq A\times A$ with $|G|\geq\delta^{\epsilon}|A|^{2}$, the $\delta$-covering numbers satisfy
$\mathcal{E}_{\delta}(A+_{G}A)+\mathcal{E}_{\delta}(\alpha\cdot A^{\wedge 2}+_{G}\beta\cdot A^{\wedge 2})\gg_{\alpha,\beta,s}C^{-O_{s}(1)}\delta^{-\epsilon}|A|,$
whenever $\delta$ is small enough.
@@ 1 Introduction
More precisely, we are interested in bounding the following quantity $\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}$ from below, where $\phi(x,y)=a_{1}x^{2}+a_{2}xy+a_{3}y^{2}$ is a non-degenerate rational quadratic form, i.e. $a_{2}^{2}-4a_{1}a_{3}\neq 0$, $X$ and $Y$ are independent $(s,C)$-Frostman random variables whose ranges are bounded, and $s\in(0,1)$.
Throughout this paper, the quadratic polynomial $\phi(x,y)$ is called special if there exist univariate real-analytic functions $h,a,b$ such that $\phi(x,y)=h\bigl{(}a(x)+b(y)\bigr{)}$. A direct computation shows that a special quadratic polynomial is one of the following two forms: $\phi(x,y)=\alpha x^{2}+\beta y^{2}$ and $\phi(x,y)=\gamma xy$, for $\alpha,\beta,\gamma\in\mathbb{R}^{*}$.
For $\phi(x,y)=x\cdot y$, we recall the following theorem due to Máthé and O’Regan in [12]. Let $X$ and $Y$ be independent $(s,C)$-Frostman random variables whose ranges are contained in $c_{1},c_{1}]$, with $0<s<1$, and $C,c_{1}\geq 1$. Then there exists a positive constant $\epsilon=\epsilon(s)$ and an integer $N=N(s,C,c_{1})$ such that for all $n>N,$ we have
$\max\left\{H_{n}(X+Y),\,H_{n}(X\cdot Y)\right\}\geq n(s+\epsilon).$ (1)
To state our main theorems, we first introduce the following definitions for bivariate random variables $(X,Y)$.
Definition 1.1.
Let $s,s_{1},s_{2}\geq 0$ and $C,C_{1},C_{2}>0$.
(i) We say that the random variable $(X,Y)$ is jointly $(s_{1},s_{2},C)$-Frostman if
$P(X\in B(x,r_{1}),\,Y\in B(y,r_{2}))\leq Cr_{1}^{s_{1}}r_{2}^{s_{2}}$
for any ball $B(x,r_{1})$ and $B(y,r_{2})$.
(ii) We say that the random variable $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman if $X$ and $Y$ are $(s_{1},C_{1})$-Frostman and $(s_{2},C_{2})$-Frostman, respectively.
(iii) We say a random variable $X$ is $(s,C)$-Frostman given $Y$ if
$P(X\in B(x,r),\,Y\in E)\leq Cr^{s}P(Y\in E)$
for any ball $B(x,r)$ and Borel set $E$. Moreover, We say that the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, if $X$ is $(s_{1},C_{1})$-Frostman given $Y$, and $Y$ is $(s_{2},C_{2})$-Frostman given $X$.
These definitions are not equivalent. At a common exponent level $(s_{1},s_{2})$, we have the implications
$\boxed{\text{\parbox{116.65646pt}{\centering independent\\ and marginally Frostman\@add@centering}}}\;\Rightarrow\;\boxed{\text{conditionally Frostman}}\;\Rightarrow\;\boxed{\text{jointly Frostman}}\;\Rightarrow\;\boxed{\text{marginally Frostman}}$
In Section 7, we provide examples illustrating each level:
  • an independent pair that is marginally (s,O(1);s,O(1))(s,O(1);\,s,O(1))-Frostman for $0\leq s\leq 1$ (Example 7.1);
  • a dependent pair that is jointly $(1/4,1/4,O(1))$-Frostman, but $(1/2,O(1);1/2,O(1))$-marginally Frostman (Example 7.6);
  • a dependent pair that is conditionally $(1/4,O(1);1/4,O(1))$-Frostman, but $(1/2,1/2;O(1))$-jointly Frostman (Example 7.7).
Our first theorem addresses the case that $X$ and $Y$ are independent real random variables, $(X,Y)$ forms marginally $(s_{1},C_{1};S_{2},C_{2})$-Frostman, and $\phi(x,y)$ is a non-degenerate rational quadratic form.
@@
Let $\phi(x,y)$ be a non-degenerate rational quadratic form, $0<s<1$ and $C_{1},C_{2},c>0$. There exist a positive constant $\epsilon=\epsilon(\phi,s)$ and an integer $N=N(\phi,s,C_{1},C_{2},c)$ such that the following statement holds: Let $X$ and $Y$ be independent real random variables whose ranges are contained in [c,c][-c,c]. Suppose that $(X,Y)$ is marginally $(s_{1},C_{1};S_{2},C_{2})$-Frostman, with $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then
$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon)$ (2)
for all $n>N$.
Here, a constant depends on $\phi$ means that it depends on the coefficients $a_{1}$, $a_{2}$ and $a_{3}$.
To show the sharpness of this theorem, we summarize the main statements of several examples in Section 7:
@@
    Any $(s,C)$-Frostman random variable $Z$ has the property that $H_{n}(Z)\geq ns-\log C+O(1)$ (see Example 7.1 for an example matching this bound). The lower bound in (2) says that the discretized entropies of $X+Y$ and $\phi(X,Y)$ cannot be small simultaneously. We show in Examples 7.1 and 7.2 that at least one of the terms $H_{n}(X+Y)$ or $H_{n}(\phi(X,Y))$ can be as small as $(1+o(1))ns$.
@@
    The positive $\epsilon$ might not exist if one of the conditions of the theorem is not satisfied, including when $s=0$ or 11 (see Example 7.1), the Frostman constant $C$ grows with $n$ (see Example 7.4), or the polynomial is degenerate (see Example 7.3).
@@
    In the statement of Theorem 1.2, the fact that NN depends on $C$ , and the condition that $X$ and $Y$ are independent are both essential (see Examples 7.4 and 7.5, respectively).
@@
    For any $0<s<1$ and any sufficiently large $n\in\mathbb{N}$, there exists an $(s,O(1))$-Frostman random variable $X$ with range contained in $[0,1]$, such that
@@
It is worth noting that the proof of Theorem 1.2 can be extended to more general settings. In this version, we present two results for dependent random variables and special polynomials. Further results will be included in a forthcoming update.
Theorem 1.3.
Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$ and $C_{1},C_{2},c>0$. There exist a positive constant $\epsilon=\epsilon(s)$ and an integer $N=N(\alpha,\beta,s,C_{1},C_{2},c)$ such that the following statement holds: Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$. Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then
$\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}\geq n(s+\epsilon),$
and
$\max\left\{\frac{H_{n}(X+Y)+H_{n}(X-Y)}{2},\,H_{n}(XY)\right\}\geq n(s+\epsilon)$
for all $n>N$.
Both Theorems 1.2 and 1.3 are consequences of the following theorem, which only requires a jointly Frostman condition.
Theorem 1.4.
Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$, $C\geq 1$ and $c>0$. There exist a positive constant $\epsilon=\epsilon(s)$ and an integer $N=N(\alpha,\beta,s,C,c)$ such that the following statement holds: For any bivariate real random variable $(X,Y)$ that is jointly $(s_{1},s_{2},C)$-Frostman and has range contained in $c,c]^{2}$, where $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$, we have
$\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}\geq n(s+\epsilon)$
for all $n>N$.
Example 7.6 will illustrate that the first term $\frac{H_{n}(X)+H_{n}(Y)}{2}$ in above formula cannot be removed, and cannot be replaced by $\frac{1}{2}H(X,Y)$ as in Theorem 1.3.
In the following, we discuss some applications of the entropy theorems.
Covering number: non-degenerate quadratic polynomials.
@@
Let $\phi(x,y)$ be a non-degenerate rational quadratic form. For $0<s<1$, there exists a positive constant $\epsilon=\epsilon(s{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},\phi})$ such that for every Borel set $A\subset{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}[0,1]}$ with $H^{s}(A)>0$, we have
$\max\left\{\underline{\dim}_{B}(A+A),\overline{\dim}_{B}(\phi(A,A))\right\}\geq s+\epsilon,$
and
$\max\left\{\overline{\dim}_{B}(A+A),\underline{\dim}_{B}(\phi(A,A))\right\}\geq s+\epsilon.$
@@
Let $\phi(x,y)$ be a non-degenerate rational quadratic form. Given $0<s<1$ and $C\geq 1$. Then there exist positive constants $\epsilon=\epsilon(\phi,s)$ such that the following holds for all $\delta>0$. Let $A\subset[0,1]$ be any $\delta$-separated set of cardinality $\lceil\delta^{-s}\rceil$. Suppose that for all intervals $I$ of length $l$ at least $\delta$, we have the non-concentration estimate
$(A\cap I)\leq Cl^{s}(\#A).$
@@
$\mathcal{E}_{\delta}(A+A)+\mathcal{E}_{\delta}(\phi(A,A))\gg_{\phi,s}C^{-O_{\phi,s}(1)}\delta^{-\epsilon}(\#A).$
These two applications are of independent interest, particularly when $\phi(x,y)$ is a non-bilinear form, and might be considered as an application of the Falconer distance problem. If $\phi(x,y)=xy$, then stronger results have been obtained by using a variety of techniques, see [5, 11, 12, 17] and references therein for current progress. Given this line of research, it is reasonable to expect different proofs of Theorem 1.5 with other methods. In the prime field setting, such a result has been proved by using tools from incidence geometry, see [9, Theorem 1.5] and [15, Theorem 1.3], for example. When $\phi(x,y)=x^{2}+y^{2}$, after posting this paper to Arxiv, Tuomas Orponen pointed out that one can use his proof of [22, Corollary 1.3] and recent results in [1] or [23] to obtain a quantitative bound for this special case.
Fractional covering number: special quadratic polynomials.
For a set $G\subseteq A\times A$, we denote
$A+_{G}A=\{x+y:\,(x,y)\in G\},\qquad\alpha\cdot A^{\wedge 2}+_{G}\beta\cdot A^{\wedge 2}=\{\alpha x^{2}+\beta y^{2}:\,(x,y)\in G\}.$
The next theorem presents an application of Theorem 1.4 on fractional covering numbers.
Theorem 1.7.
Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$ and $C\geq 1$. Then there exist positive constants $\epsilon=\epsilon(s)$ and $\delta_{0}=\delta_{0}(\alpha,\beta,s,C)$ such that the following holds for all $0<\delta\leq\delta_{0}$. Let $A\subset{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}[0,1]}$ be any $\delta$-separated set of cardinality $\lceil\delta^{-s}\rceil$. Suppose that for all intervals $I$ of length $l$ at least $\delta$, we have the non-concentration estimate
$(A\cap I)\leq Cl^{s}(\#A).$
Then, for all $G\subset A\times A$ with $|G|\geq\delta^{\epsilon}|A|^{2}$, then
$\mathcal{E}_{\delta}(A+_{G}A)+\mathcal{E}_{\delta}(\alpha\cdot A^{\wedge 2}+_{G}\beta\cdot A^{\wedge 2})\gg_{\alpha,\beta,s}C^{-O_{s}(1)}\delta^{-\epsilon}(\#A).$
Sketch the main ideas in the proof of Theorem 1.2.
Our base case is corresponding to the diagonal forms $\phi(X,Y)=\alpha X^{2}+\beta Y^{2}$ (Lemma 4.1), with independent $(s,C)$-Frostman random variables $X,Y$. Roughly speaking, by applying the submodularity formula (Lemma 2.17) on certain algebraic relations between the sum function $X+Y$ and the distance function $(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}$, we are able to reduce the problem to bounding the entropy of the distance function from below. In particular, let $W$ be an $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subset\mathbb{R}^{2}$, where $0<s<1$, the constant $C\geq 1$ is bounded and does not depend on $n$, and $R>0$. Let $W^{\prime}$ be a random variable which is i.i.d. with $W$. Then there exists a constant $\epsilon_{0}=\epsilon_{0}(s)>0$ such that
@@
for all $n\geq N_{0}$, where $N_{0}=N_{0}(s,C,R)$. Note that if the Frostman constant $C$ depends on $n$, then the inequality (3) might not be true (see Example 7.5). To prove (3), we make use of results on the Falconer distance problem. While classical estimates taken from [13, Section 4.2] are sufficient in the case $s>1/2$, we invoke the recent work of Shmerkin and Wang [24] for $s\leq 1/2$.
[…]
@@
Large dimension: Let $W$ be an $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subset\mathbb{R}^{2}$, where $1/2<s<1$, $C\geq 1$ and $R>0$. Let $W^{\prime}$ be a random variable which is i.i.d with $W$. Then for any number $t$ satisfying $t\leq 1$ and s<t<2s1/2s<t<2s-1/2, we have that for all $n\in\mathbb{N},$
@@
Small dimension: Let $X^{\prime\prime}$, $Y^{\prime\prime}$ be independent random variables with ranges contained in [c,c][-c,c], with $c>0$. Suppose that $X^{\prime\prime}$ and $Y^{\prime\prime}$ are $(s,C_{1})$-Frostman and $(s,C_{2})$-Frostman, respectively, where $0<s\leq 1/2$ and $C_{1},C_{2}\geq 1$. Let $(X^{\prime\prime\prime},Y^{\prime\prime\prime})$ be i.i.d. with $(X^{\prime\prime},Y^{\prime\prime})$. Then for any $t$ with $s<t<\frac{s+\varphi(2s)}{2}$, for all $n\in\mathbb{N}$, we have
@@
Plugging these inequalities into our method, the theorem follows.